OpenStudy (anonymous):

I will give medal and fan if you can help me :) find the approximate side length of a square board game with an area of 106 in^2

OpenStudy (anonymous):

\[area = (side)^2\] area is given. so just plug the value and find its square root

OpenStudy (anonymous):

?? how do I plug in something that isnt there :(

OpenStudy (anonymous):

area = 106 \[106 = (side)^2\] now take square root

OpenStudy (anonymous):

of 106?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont know how to find the square root the correct way anymore it has been forever I used to be a pro at it by using the upside down cake

OpenStudy (anonymous):

in like 9th grade im in 12th now and just dont remember

OpenStudy (anonymous):

\[\sqrt{106}\] find its factor that is 53 and 2 but both are prime so it will \[\sqrt{53*2}\] thats your side or you can write it \[\sqrt{106}\]

OpenStudy (anonymous):

 this is what i did